# Parametric pair production of collective excitations in a Bose-Einstein condensate

### v1

@@ Abstract
By exciting the transverse breathing mode of an elongated Bose-Einstein condensate, we parametrically produce longitudinal collective excitations in a pairwise manner. This process also referred to as Faraday wave generation, can be seen as an analog to cosmological particle production. Building upon single particle detection, we investigate the early time dynamics of the exponential growth and compare our observation with a Bogoliubov description. The growth rate we observe experimentally is in very good agreement with theoretical predictions, demonstrating the validity of the Bogoliubov description and thereby confirming the smallness of quasiparticle interactions in such an elongated gas. We also discuss the presence of oscillations in the atom number, which are due to pair correlations and to the rate at which interactions are switched off.
@@ II Experimental setup and model
[…]
[width=5.51in]fig1.pdf
FIG. 1: Experimental parametric excitation of quasiparticles. a, Diagram of the experimental setup. b, By modulating dipole trap laser power $\mathcal{P}_{las}$, the breathing mode of the BEC is excited. c, The BEC width $\sigma$ oscillates as long as the cloud is kept in the trap for a duration $\Delta t$. The breathing mode parametrically excites two longitudinal excitations with opposite momenta. d, Time evolution of the longitudinal density profile of the atom velocity distribution. The dark band at $v_{z}=0$ is the condensate. After a hold time of about 2 ms, sidebands appear on either side of the condensate.
Modulation protocol – We produce a cigar-shaped BEC with typically 15,000 atoms and at a temperature between 30 and 45 nK, in a crossed dipole trap with longitudinal and transverse frequencies of 40 Hz and 1 kHz, respectively. The experimental setup and procedure, sketched in Fig. 1a, are the same as in Ref. [17], with the main difference being the number of atoms in the BEC and its temperature, which we control by adjusting the final trap power. To excite the gas, the power of the vertical trapping laser is modulated at twice the transverse trap frequency $\omega_{\perp}$, for four oscillation periods, see Fig. 1b. This modulation excites the transverse breathing mode of the BEC: its width $\sigma$ continues to oscillate with the same amplitude for an additional hold time $\Delta t$ after the modulation ends [56], as drawn in Fig. 1c (see also the experimental data in Fig. 3a). This breathing mode is not damped by thermal phonons [57] but rather couples to longitudinal collective excitations i.e. Faraday waves [58].
Detecting atoms – When the trap is switched off, density decreases and atoms that carry a collective excitation separate from the BEC, forming two sidebands, as shown in Fig. 1d. Atoms fall onto a microchannel plate detector located 46 cm below the trap, which records the arrival time and position of single atoms with a quantum efficiency of 25(5)% [59]. The time of flight is sufficiently long that the arrival time and position at the detector accurately reflect the three-dimensional velocity of the atoms at the moment the trap and interactions were switched off. The detector is protected from the vertical trapping laser; thus the 13 µs Raman pulse that transfers the atoms to a magnetically insensitive state after the trap switch off also imparts a transverse momentum, ensuring that the atoms are detected on the unshielded region of the detector. In some data sets, a 1 ms velocity-selective Bragg pulse deflects 97% of the BEC atoms upwards at 48 mm/s, while not affecting the sidebands [60]. The Bragg pulse helps prevent saturation of the detector in the vicinity of the sidebands, which may alter their shape and intensity. However, in the data shown in this work the presence of the Bragg pulse had no observable effect on the sidebands.
@@
Here, $\hbar$ is the reduced Planck constant, $m$ is the mass of the atoms, $n_{1}$ is the one-dimensional density and $\dot{\omega}_{k}$ represents the time derivative of $\omega_{k}$. The effective 1D interaction strength $g_{1}$ is proportional to the mean transverse density, and therefore inversely proportional to the cross-sectional area of the BEC. Assuming a Gaussian transverse profile and extending the result of Refs [63, 64] to the anisotropic case, we find $g_{1}=2\hbar^{2}a_{s}/(m\sigma_{x}\sigma_{y})$, where $a_{s}$ is the atomic $s$-wave scattering length and where $x$ and $y$ correspond to the principal axes along which the transverse profile has minimal and maximal width. During the breathing oscillation, $\sigma_{x}$ and $\sigma_{y}$ vary in time which in turn modulates the dispersion relation in Eq. (1). In the absence of oscillation, $\dot{\omega}_{k}=0$ and the second term in Eq. (1) vanishes. Thus, each quasiparticle mode $\hat{b}_{k}$ evolves independently as $\hat{b}_{k}\propto e^{i\omega_{k}t}$. However, when $g_{1}$ varies, we have $\dot{\omega}_{k}\neq 0$ and the evolution of modes with opposite wavevector are coupled. In particular, when $\dot{\omega}_{k}$ oscillates at $2\omega_{\perp}$, the two modes with $\omega_{\pm k}=\omega_{\perp}$ are parametrically excited and squeezed. The number of quasiparticles $\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle$ increases exponentially, as does their anomalous correlation $|\langle\hat{b}_{k}\hat{b}_{-k}\rangle|$, signaling a pair production process. The rate of the exponential growth depends on the breathing amplitude, which itself depends on the amplitude and duration of the laser power modulation.
@@ III Exponential creation of quasiparticles
[…]
[width=]fig2.pdf
FIG. 2: Detected atom number in each sideband as a function of the hold time. In this data set the laser power was modulated by ±15%. The atom number was counted in each voxel as defined in the text. The data was fitted to Eq. (2) in the exponential regime, that is between 0.5 and 4 ms. The solid blue and dashed orange lines show the fits for the negative and positive velocity sidebands respectively. The fits give a reduced chi-squared of $\chi_{\nu}^{2}=1.0$ and $\chi_{\nu}^{2}=1.2$ respectively.
@@
To quantify the growth, we define a 3D volume in velocity space, a voxel, whose longitudinal size is roughly that of one mode [17]. For each hold time $\Delta t$, the voxel is centered at the velocity for which the atom density is maximum. In Fig. 2, the evolution of the number of atoms in the two sidebands is plotted as a function of hold time, confirming the exponential growth. The oscillations shown here have also been observed in other experiments [10, 19, 16, 13] and we will discuss their interpretation in section V.
@@
The model fits well to the data in Fig. 2. The physical meaning of some fit parameters, whose theoretical expression is given in the appendix, is as follows. The offset $n_{0}$ accounts for other non-excited modes which are detected in the voxel analysis. The exponential growth is multiplied by an overall prefactor $\Delta n$, which characterizes the fluctuations that seed the parametric growth. The parameter $G_{k}$ characterizes the growth and $A_{k}$ the amplitude of the oscillations. In our fit, $n_{0}$ and $\Delta n$ are sensitive to the analysis voxel size, while $G_{k}$ and $A_{k}$ are robust. These will be discussed in the next sections.
@@ IV Measurement of the growth rate
where $a$ is the modulation amplitude and $\omega_{d}$ the driving frequency, the growth dynamics is analytical. Without damping, the growth rate for the resonant mode $\omega_{k}=\omega_{d}/2$ in a homogeneous gas is predicted to be [36]
@@
[…]
[width=]fig3.pdf
FIG. 3: Comparison between the experimental and homogeneous undamped theoretical rate. a, The asymmetric breathing mode of the BEC is fitted with a sine function along both axes. The modulation parameters are the same as in Fig. 2. b, The relative, effective 1D interaction strength. c, The fitted growth rate for different modulation amplitudes on the vertical axis is compared to the prediction of Eq. (4) on the horizontal axis. The theoretical growth rate is also shown as a line of unit slope.
@@
Results – We show in Fig. 3c the measured growth rate $G_{k}^{\rm exp}$ as a function of the theoretical growth rate $G_{k}^{\rm th}$, evaluated using Eq. (4). The line of unit slope thus corresponds to the theoretical prediction, and the measured growth rates are seen to be in very good agreement with it the predicted ones. Recently, it was shown that the growth rate should be sensitive to interactions between quasiparticles [61] arising from higher-order terms in the Hamiltonian, which are not taken into account within the Bogoliubov description. The dominant effect is the presence of a decay rate that acts counter to the growth of quasiparticle populations and correlations [61, 67]. With our experimental parameters, the magnitude of this decay is predicted to be on the order of 0.1 ms-1, and while our data are not sufficiently precise to confirm such a small reduction, they do confirm that any decay rate present cannot be much larger.
@@ V Microscopic interpretation and scaling of the oscillations
Bogoliubov transformation – During the parametric excitation, the number of quasiparticles grows exponentially. We refer here to the quasiparticle operators as the operators $\hat{b}_{k}$ which diagonalize the Hamiltonian at any time. These quasiparticles correspond to superpositions of $N$ atoms moving in one direction and N1N-1 moving in the other. Since quasiparticles are produced coherently in $\pm k$ pairs, the superposition of these moving atoms creates a standing wave in the atomic density, which oscillates at frequency $\omega_{k}$ as observed in Refs. [10, 12, 11, 19]. This standing wave corresponds to a superposition of equal numbers of atoms at momenta $k$ and k-k, whose number is minimal whenever the density is momentarily stationary. As this occurs twice per cycle, it oscillates at frequency $2\omega_{k}$. At (almost) any instant there is thus an excess of moving atoms above the minimum, and this excess can be interpreted physically as driving the movement of the density pattern, i.e. as those atoms which are moving from regions of decreasing density into regions of increasing density.
@@
which is approximately equal to 0.7 with our parameters. However, experiments that measure atoms in momentum space do so either after time-of-flight expansion (as in our case) or by imaging atoms following a phase-space rotation [69, 70], as in Ref. [16]. In both cases, the measurement is performed once the atoms cease to interact and quasiparticles and atom operators are formally identical: $u_{k}=1$ and $v_{k}=0$. Then, the amplitude of the observed oscillation depends on the mapping of the collective excitation state to the atomic one and this depends on the rate of change of the quasiparticle frequency $\dot{\omega}_{k}/\omega_{k}$. This can be seen from Eq. (1) where a high value of $\dot{\omega}_{k}/\omega_{k}$ induces a mixing between the $k$ and k-k. This point has also been discussed in the context of other analog gravity experiments in Ref. [71].
@@
[…]
[width=]fig4.pdf
FIG. 4: A more adiabatic mapping of the quasiparticles to the atoms. a, Numerical solution and b, experimental plot of the evolution of the number of atoms when the transverse laser power, shown in the inset, is abruptly turned off (solid and round markers, in red) or ramped down in 1 ms (dashed line and square markers, in green).
Intermediate case – In our experiment, the interaction strength is driven by the transverse width of the BEC $g_{1}\propto\sigma^{-2}$. Even if the dipole trap is turned off instantaneously, the transverse width increases on a time scale of the order of $\omega_{\perp}^{-1}$. As the frequency of the excited quasiparticles corresponds precisely to this timescale, we are neither in the adiabatic nor in the diabatic regime. The observed oscillation amplitudes in Figs. 2 and 4b are about 0.5, smaller than $A_{k}^{\rm inst}=0.7$ and close to the amplitude found numerically (see Fig. 4a). Although we cannot strictly reach it experimentally, we can approach the adiabatic case by slowly ramping down the transverse laser power. The resulting evolution of the atom number for a 1 ms ramp is shown in green squares Fig. 4b. One sees that the oscillations, if present, are quite weak. On the other hand, the total number of atoms is higher than for the sudden turnoff. This is because the pair production mechanism continues to operate during the ramping down of the laser power (see Appendix A.4).
@@ VI Conclusion
Our observations of exponential growth and oscillation of the particle number are in agreement with Bogoliubov theory as applied in Ref. [37]. The data clearly shows the parametric nature of the process and the fact that the excitations are generated in a pairwise manner.
The measured growth rate is in very good agreement with the predicted value, even in the presence of a trap inhomogeneity. The precision of our measurements cannot isolate and measure the small effect of quasiparticle interactions, which should reduce the growth rate [61]. Nevertheless, our results do confirm the smallness of this decay, if present. Future work will aim to improve the experimental procedure in order to increase the signal-to-noise ratio and further isolate this effect. In addition, we envision repeating the above experiments in a square potential [73] in order to realize a situation closer to the present theoretical model. Alternatively, the model could be improved by including the density inhomogeneities. Also, exciting the gas with a Feshbach resonance would allow to excite other $k$ modes while our excitation method is restricted to modes satisfying $\omega_{k}=\omega_{\perp}$  [^3] By varying the excitation frequency of the laser power, we could in principle excite other modes [40]. However, the procedure did not yield such a clear parametric growth as in Fig. 2.. This perspective is interesting because the predicted decay rate of Ref. [61] depends on $k$.

### v2

@@ Abstract
By exciting the transverse breathing mode of an elongated Bose-Einstein condensate, we parametrically produce longitudinal collective excitations in a pairwise manner. This process also referred to as Faraday wave generation, can be seen as an analog to cosmological particle production. Building upon single particle detection, we investigate the early time dynamics of the exponential growth and compare our observations with a Bogoliubov description. The growth rate we observe experimentally is in very good agreement with theoretical predictions, demonstrating the validity of the Bogoliubov description and thereby confirming the smallness of quasiparticle interactions in such an elongated gas. We also discuss the presence of oscillations in the atom number, which are due to pair correlations and to the rate at which interactions are switched off.
@@ II Experimental setup and model
Figure 1: Experimental parametric excitation of quasiparticles. a, Diagram of the experimental setup. b, By modulating dipole trap laser power $\mathcal{P}_{las}$, the breathing mode of the BEC is excited. c, The BEC width $\sigma$ oscillates as long as the cloud is kept in the trap for a duration $\Delta t$. The breathing mode parametrically excites two longitudinal excitations with opposite momenta. d, Time evolution of the longitudinal density profile of the atom velocity distribution. The dark band at $v_{z}=0$ is the condensate. After a hold time of about 2 ms, sidebands appear on either side of the condensate.
Modulation protocol – We produce a cigar-shaped BEC with typically 15,000 atoms and at a temperature between 30 and 45 nK, in a crossed dipole trap with longitudinal and transverse frequencies of 40 Hz and 1 kHz, respectively. The experimental setup and procedure, sketched in Fig. 1a, are the same as in Ref. [17], with the main difference being the number of atoms in the BEC and its temperature, which we control by adjusting the final trap power. To excite the gas, the power of the vertical laser $P_{\rm las}$, which confines the gas transversely, is modulated at twice the transverse trap frequency $\omega_{\perp}$, for four oscillation periods, see Fig. 1b. This modulation excites the transverse breathing mode of the BEC: its width $\sigma$ continues to oscillate with the same amplitude for an additional hold time $\Delta t$ after the modulation ends [56], as drawn in Fig. 1c (see also the experimental data in Fig. 3a). This breathing mode is not damped by thermal phonons [57] but rather couples to longitudinal collective excitations i.e. Faraday waves [58].
Detecting atoms – When the trap is switched off, the density decreases and atoms that carry a collective excitation separate from the BEC, forming two sidebands, as shown in Fig. 1d. Atoms fall onto a microchannel plate detector located 46 cm below the trap, which records the arrival time and position of single atoms with a quantum efficiency of 25(5)% [59]. The time of flight is sufficiently long that the arrival time and position at the detector accurately reflect the three-dimensional velocity of the atoms at the moment the trap and interactions were switched off. The detector is protected from the vertical trapping laser; thus the 13 µs Raman pulse that transfers the atoms to a magnetically insensitive state after the trap switch off also imparts a transverse momentum, ensuring that the atoms are detected on the unshielded region of the detector. In addition, in some data sets, a 1 ms velocity-selective Bragg pulse deflects 97% of the BEC atoms upwards at 48 mm/s, while not affecting the sidebands [60]. The Bragg pulse helps prevent saturation of the detector in the vicinity of the sidebands, which may alter their shape and intensity.
@@
Here, $\hbar$ is the reduced Planck constant, $m$ is the mass of the atoms, $n_{1}$ is the one-dimensional density and $\dot{\omega}_{k}$ represents the time derivative of $\omega_{k}$. The effective 1D interaction strength $g_{1}$ is proportional to the mean transverse density, and therefore inversely proportional to the cross-sectional area of the BEC. Assuming a Gaussian transverse profile and extending the result of Refs [63, 64] to the anisotropic case, we find $g_{1}=2\hbar^{2}a_{s}/(m\sigma_{x}\sigma_{y})$, where $a_{s}$ is the atomic $s$-wave scattering length and where $x$ and $y$ correspond to the principal axes along which the transverse profile has minimal and maximal width. During the breathing oscillation, $\sigma_{x}$ and $\sigma_{y}$ vary in time which in turn modulates the dispersion relation in Eq. (1). In the absence of oscillation, $\dot{\omega}_{k}=0$ and the second term in Eq. (1) vanishes. Thus, each quasiparticle mode $\hat{b}_{k}$ evolves independently as $\hat{b}_{k}\propto e^{-i\omega_{k}t}$. However, when $g_{1}$ varies, we have $\dot{\omega}_{k}\neq 0$ and the evolution of modes with opposite wavevector are coupled. In particular, when $\dot{\omega}_{k}$ oscillates at $2\omega_{\perp}$, the two modes with $\omega_{\pm k}=\omega_{\perp}$ are parametrically excited and squeezed. The number of quasiparticles $\langle\hat{b}_{k}^{\dagger}\hat{b}_{k}\rangle$ increases exponentially, as does their anomalous correlation $|\langle\hat{b}_{k}\hat{b}_{-k}\rangle|$, signaling a pair production process. The rate of the exponential growth depends on the breathing amplitude, which itself depends on the amplitude and duration of the laser power modulation.
@@ III Exponential creation of quasiparticles
Figure 2: Detected atom number in each sideband as a function of the hold time. In this data set the laser power was modulated by ±15%. The atom number was counted in each voxel as defined in the text. The data was fitted to Eq. (2) in the exponential regime, that is between 0.5 and 4 ms. The solid blue and dashed orange lines show the fits for the negative and positive velocity sidebands respectively. The fits give a reduced chi-squared of $\chi_{\nu}^{2}=1.0$ and $\chi_{\nu}^{2}=1.2$ respectively.
@@
To quantify the growth, we define a 3D volume in velocity space, a voxel, whose longitudinal size is roughly that of one mode [17]. For each hold time $\Delta t$, each voxel is centered at the velocity for which the atom density is maximum, one on the positive velocity sideband and the other one on the negative one. In Fig. 2, the evolution of the number of atoms in the two voxels is plotted as a function of hold time, confirming the exponential growth. The oscillations shown here have also been observed in other experiments [10, 19, 16, 13] and we will discuss their interpretation in section V.
@@
The model fits well to the data in Fig. 2. The physical meaning of some fit parameters, whose theoretical expression is given in the appendix, is as follows. The offset $n_{0}$ accounts for other non-excited modes which are detected in the voxel analysis. The exponential growth is multiplied by an overall prefactor $\Delta n$, which characterizes the fluctuations that seed the parametric growth. The parameter $G_{k}$ characterizes the growth and $A_{k}$ the amplitude of the oscillations. In our fit, $n_{0}$ and $\Delta n$ are sensitive to the analysis voxel size, while $G_{k}$ and $A_{k}$ are not. These will be discussed in the next sections.
@@ IV Measurement of the growth rate
where $a$ is the modulation amplitude and $\omega_{d}$ the driving frequency, the growth dynamics can be solved analytically. Without damping, the growth rate for the resonant mode $\omega_{k}=\omega_{d}/2$ in a homogeneous gas is predicted to be [36]
@@
Figure 3: Comparison between the experimental and homogeneous undamped theoretical rate. a, The asymmetric breathing mode of the BEC is fitted with a sine function along both axes. The modulation parameters are the same as in Fig. 2. b, The relative, effective 1D interaction strength. c, The fitted growth rate for different modulation amplitudes on the vertical axis is compared to the prediction of Eq. (4) on the horizontal axis. The theoretical growth rate is also shown as a line of unit slope.
@@
Results – We show in Fig. 3c the measured growth rate $G_{k}$ as a function of the theoretical growth rate $G_{k}^{\rm th}$, evaluated using Eq. (4). The line of unit slope thus corresponds to the theoretical prediction, and the measured growth rates are seen to be in very good agreement with it the predicted ones. Recently, it was shown that the growth rate should be sensitive to interactions between quasiparticles [61] arising from higher-order terms in the Hamiltonian, which are not taken into account within the Bogoliubov description. The dominant effect is the presence of a decay rate that acts counter to the growth of quasiparticle populations and correlations [61, 67]. With our experimental parameters, the magnitude of this decay is predicted to be on the order of 0.1 ms-1, and while our data are not sufficiently precise to confirm such a small reduction, they do confirm that any decay rate present cannot be much larger.
@@ V Microscopic interpretation and scaling of the oscillations
Bogoliubov transformation – During the parametric excitation, the number of quasiparticles grows exponentially. We refer here to the quasiparticle operators as the operators $\hat{b}_{k}$ which diagonalize the Hamiltonian at any time. These quasiparticles correspond to superpositions of $N$ atoms moving in one direction and N1N-1 moving in the other. Since quasiparticles are produced coherently in $\pm k$ pairs, the superposition of these moving atoms creates a standing wave in the atomic density, which oscillates at frequency $\omega_{k}$ as observed in Refs. [10, 12, 11, 19]. This standing wave corresponds to a superposition of equal numbers of atoms at momenta $k$ and k-k, whose number is minimal whenever the density is momentarily stationary. As this occurs twice per cycle, the atom numbers oscillate at frequency $2\omega_{k}$. At (almost) any instant there is thus an excess of moving atoms above the minimum, and this excess can be interpreted physically as driving the movement of the density pattern, i.e. as those atoms which are moving from regions of decreasing density into regions of increasing density.
@@
which is approximately equal to 0.7 with our parameters. However, experiments that measure atoms in momentum space do so either after time-of-flight expansion (as in our case) or by imaging atoms following a phase-space rotation [69, 70], as in [16]. In both cases, the measurement is performed once the atoms cease to interact and quasiparticles and atom operators are formally identical: $u_{k}=1$ and $v_{k}=0$ in Eq. (5). The amplitude of the observed oscillation in $n^{\rm at}_{k}$ then depends on the details of the measurement protocol which modifies the state of the collective excitations, e.g. the values $\langle{b}_{k}^{\dagger}{b}_{k}\rangle$ and $\langle{b}_{-k}{b}_{k}\rangle$, while mapping it to that of the atoms. For a time-of-flight expansion, this modification depends on the rate of change of the quasiparticle frequency $\dot{\omega}_{k}/\omega_{k}$ as can be seen from Eq. (1) where a high value of $\dot{\omega}_{k}/\omega_{k}$ induces a mixing between the $k$ and k-k modes, see Appendix A.2 for details. This point has also been discussed in the context of other analog gravity experiments in [71].
@@
Figure 4: A more adiabatic mapping of the quasiparticles to the atoms. a, Numerical solution and b, experimental plot of the evolution of the number of atoms when the transverse laser power, shown in the inset, is abruptly turned off (solid and round markers, in red) or ramped down in 1.5 ms (dashed line and square markers, in green).
Intermediate case – In our experiment, the interaction strength is driven by the transverse width of the BEC $g_{1}\propto\sigma^{-2}$. Even if the dipole trap is turned off instantaneously, the transverse width increases on a time scale of the order of $\omega_{\perp}^{-1}$. As the frequency of the excited quasiparticles corresponds precisely to this timescale, we are neither in the adiabatic nor in the diabatic regime. The observed oscillation amplitudes in Figs. 2 and 4b are about 0.5, smaller than $A_{k}^{\rm inst}=0.7$ and close to the amplitude found numerically (see Fig. 4a). Although we cannot strictly reach it experimentally, we can approach the adiabatic case by slowly ramping down the transverse laser power. The resulting evolution of the atom number for a 1.5 ms ramp is shown in green squares Fig. 4b. One sees that the oscillations, if present, are quite weak. On the other hand, the total number of atoms is higher than for the sudden turnoff. This is because the pair production mechanism continues to operate during the ramping down of the laser power (see Appendix A.4).
@@ VI Conclusion
In this paper, we have studied the production of quasiparticles in a parametrically driven BEC. Our observations of exponential growth and oscillation of the particle number are in agreement with Bogoliubov theory as applied in Ref. [37]. The data clearly shows the parametric nature of the process and the fact that the excitations are generated in a pairwise manner.
The measured growth rate is in very good agreement with the predicted value, even in the presence of a trap inhomogeneity. The precision of our measurements cannot isolate and measure the small effect of quasiparticle interactions, which should reduce the growth rate [61]. Nevertheless, our results do confirm the smallness of this decay, if present. Future work will aim to improve the experimental procedure in order to increase the signal-to-noise ratio and further isolate this effect. Higher experimental precision may require a closer correspondence between the experiment and the model, especially if we wish to compare the value of the damping of the collective excitations with the prediction of Ref. [61]. Thus we envision repeating the above experiments in a square potential [73]. Alternatively, the model could be improved by including the density inhomogeneities. Also, exciting the gas with a Feshbach resonance would allow to excite other $k$ modes while our excitation method is restricted to modes satisfying $\omega_{k}=\omega_{\perp}$  [^3] By varying the excitation frequency of the laser power, we could in principle excite other modes [40]. However, the procedure did not yield such a clear parametric growth as in Fig. 2.. This perspective is interesting because the predicted decay rate of Ref. [61] depends on $k$.